Redshift measurements show that a galaxy is moving away from the Earth at the speed of 60000 km/s. Taking the Hubble constant H, as 2.2 x 10-18s¹ and 1 light-year as 10¹3 km, calculate and show your working for:


a)how far the galaxy is from the Earth distance


b)how long it has taken light from the galaxy to reach the Earth


c) the time the galaxy and the Milky Way have been travelling apart


d) Explain the significance of the time you calculated in c part of the question


1
Expert's answer
2022-05-31T16:11:25-0400

"r=\frac vH=27.3\cdot 10^{21}~km=2.73\cdot 10^9~\text{l.y},"

"t=\frac rc=2.79\cdot 10^9~\text{years},"

"t_0=\frac t 2\sqrt{1-\frac{v^2}{c^2}}=1.34\cdot 10^9~\text{years}."